On a planet that has no atmosphere, a rocket 15.8 m tall is resting on its launch pad. Freefall acceleration on the planet is 4.45 m/s2. A ball is dropped from the top of the rocket with zero initial velocity. (a) How long does it take to reach the launch pad? (b) What is the speed of the ball just before it reaches the ground?
1
Expert's answer
2010-11-14T17:38:16-0500
From the equations of the motion
t = √(2h/a) = √(2*15.8/4.45) = 2.66 s.
v = v0 + at = 0 + 4.45*2.66 = 11.84 m/s.